OpenStudy (lapshaman):

how to solve, math test tomorrow, plz help |x-8|=4 thanks!

OpenStudy (littlebird):

http://www.tiger-algebra.com/drill/abs%28x-8%29=4/

OpenStudy (lapshaman):

thanks

OpenStudy (lapshaman):

it should only have one solution

Nnesha (nnesha):

so you're going to use that calculator tomorrow ?

OpenStudy (lapshaman):

what calculator?

Nnesha (nnesha):

tiger website or whatever that is

Nnesha (nnesha):

lets start solving for x

OpenStudy (littlebird):

the website came with an explanation

OpenStudy (lapshaman):

no, it doesn't give the correct answer

OpenStudy (littlebird):

woops. I didnt really read that thing thoroughly

Nnesha (nnesha):

absolute value \[a=\left| c \right|\] has two solution a=c AND a=-c \[ \left| a-b \right|=c\] would be a-b=c AND a-b=-c

OpenStudy (littlebird):

good thing that was caught

OpenStudy (lapshaman):

ok, cool, thanks!

Nnesha (nnesha):

when we remove the absolute bar we should get two equations a-b=c and a-b=-c and if the right side is negative then you can't solve \[\rm \left| a-b \right|=-c\] that means there is no solution

OpenStudy (lapshaman):

Ok, thanks!

Nnesha (nnesha):

do you know why we get two equations ? remember if you take absolute bar of negative number u will get positive answer I-xI= x so we don't know if the a-b is negative or positive that's why we should write it as a-b=c and a-b=-c after removing the bars

Nnesha (nnesha):

np

OpenStudy (lapshaman):

what do I do if <= is the case?

Nnesha (nnesha):

do you mean inequalities ?? if yes then that's almost the same but for negative one you should flip the sign like \[\rm \left| a-b \right| \ge c \] would be \[\rm a-b \ge c ~or~ a-b \le -c\]